# https://ssugarwh.github.io/consistentid.github.io/

### v1

@@ 3.3 Training and Inference Details
During the training process, we optimize only the parameters of the facial encoder and the projection module within the overall facial ID feature extractor, while maintaining the parameters of the pre-trained diffusion model in a frozen state. The training data for ConsistentID consists of facial image-text pairs. In ConsistentID, to enhance text controllability, we prioritize the caption as the primary prompt and concatenate it with more detailed descriptions of facial regions extracted from LLaVA1.5, forming the ultimate textual input. Regarding training loss functions, they align with those used in the original stable diffusion models and are expressed as:
@@
The total loss function is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{noise}}+\mathcal{L}_{\text{loc}}$.
During the inference process, we employ a straightforward delayed primacy condition as similar to Fastcomposer. This allows the use of a separate text representation initially, followed by enhanced text representation after a specific step, effectively balancing identity preservation and editability. More discussions are provided in the supplementary materials.
3.4 Fine-grained Human Dataset Construction
@@
To tackle this issue, we introduce a dataset pipeline outlined in the supplementary materials to create our dataset FGID. This dataset encompasses comprehensive fine-grained ID information and detailed facial descriptions, which are crucial for training the ConsistentID model.
Our FGID dataset, comprising 525,258 images is a compilation from diverse public datasets such as FFHQ [17], CelebA[27], SFHQ[2], etc, including about 524,000 training data. This dataset is tailored to encompass both whole-face and facial feature information, providing richer textual and visual details for model training [^1] In the future, we intend to introduce additional datasets, including higher-level datasets like LAION-Face[61] and self-collected multi-ID data, to further augment diversity and information content.. More details about the FGID dataset are available in the supplementary materials, providing information on its statistics, characteristics, and related ethical considerations. In the following, we elaborate on the multi-modal data handling process. Textual Data: For textual facial descriptions, MLLM LLaVA1.5 is utilized to extract detailed information using an embedded prompt ‘Please describe the people in the image, including their gender, age, clothing, facial expressions, and any other distinguishing features’. Visual Data: BiSeNet [57] and InsightFace[7] models are deployed to capture both whole-face ID information and facial feature information, ensuring comprehensive identity details.
4 Experiment
Experimental Implements. In ConsistentID, we employ the SD1.5 model as the foundational text-to-image model. For the fine-grained multimodal feature extractor, we initialize the parameters of all text encoders and image encoders with CLIP-ViT-H [44] and utilize its image projection layers to initialize the learnable projection in the overall facial ID feature extractor. The entire framework is optimized using Adam [19] on 8 NVIDIA 3090 GPUs, with a batch size of 16. We set the learning rate for all trainable modules to $1\times 10^{-4}$. During training, we probabilistically remove 50% of the background information from the characters with a 50% probability to mitigate interference. Additionally, to enhance generation performance through classifier guidance, there is a 10% chance of replacing the original updated text embedding with a zero text embedding. In inference, we employ delayed topic conditioning [17, 54] to resolve conflicts between text and ID conditions. We utilize a 50-step DDIM [45] sampler, and the scale for classifier-guided settings is set to 5.
Experimental Metrics. To assess the effectiveness and efficiency of ConsistentID, we employ six widely used metrics [40]: CLIP-I [8], CLIP-T [35], DINO [6], FaceSim [43], FID [10], and inference speed. CLIP-T measures the average cosine similarity between prompt and image CLIP embeddings, evaluating ID fidelity. CLIP-I calculates the average pairwise cosine similarity between CLIP embeddings of generated and real images, assessing prompt fidelity. DINO represents the average cosine similarity between ViT-S/16 embeddings of generated and real images, indicating fine-grained image-level ID quality. FaceSim determines facial similarity between generated and real images using FaceNet [43]. FID gauges the quality of the generated images [10]. Inference speed denotes the calculation time under the same running environment. Additionally, we introduce a novel metric, FGIS (fine-grained identity similarity), to assess ID quality at the region level. FGIS is computed as the average cosine similarity between DINO embeddings of the generated facial regions in reference and generated images. A higher FGIS value indicates increased ID fidelity in the generated facial regions.
[…]
Fastcomposer[54] 27.8 67.0 68.4 75.2 77.7 372.8 10
IP-Adapter[56] 27.6 75.0 74.5 75.6 73.4 320.0 13
Photomaker[22] 30.7 71.7 72.6 69.3 73.2 336.5 17
InstantID[53] 30.3 68.2 77.6 76.5 78.3 271.9 19
ConsistentID 31.1 76.7 78.5 77.2 81.4 312.4 16
Table 1: Quantitative comparison of the universal recontextualization setting on the MyStyle test dataset. The benchmark metrics assessed text consistency (CLIP-T), the preservation of coarse- and fine-grained ID information (CLIP-I, DINO, FaceSIM, and FGIS), generation quality (FID), and inference efficiency (speed in seconds).
4.1 Comparation Results
To demonstrate the effectiveness of ConsistentID, we conduct a comparative analysis against state-of-the-art methods, including Fastcomposer [54], IP-Adapter[56], Photomaker [22], and InstantID [53]. Our focus is on personalized generation utilizing only one reference image. We utilize the officially provided models, use the default parameters for each method, and restrict the inference to a single reference image. In alignment with the Photomaker methodology, we employ the Mystyle [31] dataset for quantitative assessment and incorporate over ten identity datasets for visualization.
Quantitative results: Following Photomaker [22], we use the test dataset from Mystyle [31], using MLLM LLaVA1.5 to obtain facial descriptions during inference. The quantitative comparison is conducted under the universal recontextualization setting, utilizing a set of metrics to benchmark various aspects.
The results are showcased in Table 1. A thorough analysis of the table demonstrates that ConsistentID consistently outperforms other methods across most evaluated metrics, and surpasses other IP-Adapter-based methods in terms of generation efficiency. This is attributed to ConsistentID’s fine-grained ID preservation capability and the efficiency of the lightweight multimodal facial prompt generator. Regarding the FID metric, the lower performance could be primarily attributed to the limited generative capability of the base model SD1.5.
Visualized comparisons under different scenarios: To visually demonstrate the advantages of ConsistentID, we present the text-edited generation results of all methods using reference images of five distinct identities in Figure 4. This visualization highlights ConsistentID’s capability to produce vibrant and realistic images, with a particular emphasis on facial features. To further elucidate this observation, we selectively magnify and compare specific facial details across all methods in four identities, as depicted in Figure 5. Our model showcases exceptional ID preservation capabilities in facial details, especially in the eyes and nose, attributed to fine-grained multimodal prompts and facial regions’ ID information.
Figure 6: Qualitative comparison of our model with other models on two special tasks: stylization and action instruction.
To validate the accurate text understanding capability, we additionally show style-based and action-based text-edited results in Figure 6. We observe that the generated images from InstantID lack sufficient flexibility in facial poses. This limitation is likely attributed to Controlnet-based prompt insertion methods, which may easily overlook textual prompts. Simultaneously, we notice that, while Photomaker can accurately comprehend textual and visual prompts, it lacks the ID consistency of facial regions. In contrast, our ConsistentID achieves optimal generation results due to its precise understanding of textual and visual prompts. This further emphasizes the significance of multimodal fine-grained ID information. To fully show the advantages of our ConsistentID, more visualized comparisons are provided in the supplementary materials, including comparative experiments with fine-tuning-based models Dreambooth [40], Textual Inversion [8], and CustomDiffusion [20].
Moreover, we compare ConsistentID with models specifically designed using IP-Adapter as the base model in Figure 7. From the figure, it is evident that this series of works currently falls short in achieving highly detailed ID preservation in facial areas without fine-grained textual and visual prompts. In contrast, ConsistentID exhibits a robust capability to preserve the integrity of facial ID and seamlessly blend it into various styles, utilizing multimodal fine-grained prompts. This comparison emphasizes the superiority of ConsistentID in retaining identity while simultaneously maintaining flexibility and control over style.
4.2 Human Study
We also investigate user preferences regarding image fidelity, fine-grained ID fidelity, and overall ID fidelity through surveys. In Figure 9, we present a visualization of the proportion of total votes received by each method. Across all three metric dimensions, ConsistentID holds the most significant share.
Figure 8: User preferences across image fidelity, fine-grained ID fidelity, overall ID fidelity for different methods
Figure 9: Ablation study of the number of facial regions.
4.3 Ablation Study
Facial ID Types: We conducted an ablation study on facial ID, considering three variations: using solely overall facial ID, incorporating our designed fine-grained ID, and concurrently leveraging both facial ID and our designed fine-grained ID. Results in Table 2 show that using fine-grained ID features alone enhances ID consistency in facial regions, indicated by an increased DINO value. However, overall ID preservation, assessed by FaceSim, cannot be adequately maintained. Utilizing only overall facial ID enhances facial similarity, but facial feature consistency diminishes, performing significantly worse in CLIP-I and DINO metrics compared to using only facial features. Combining both overall facial ID and fine-grained ID leads to a balanced improvement in the model’s overall ID fidelity compared to using only overall facial ID and fine-grained ID.
$\mathcal{L}_{\text{noise}}$ $\mathcal{L}_{\text{loc}}$ CLIP-I $\uparrow$ DINO $\uparrow$ FGIS $\uparrow$ LLaVA1.5 CLIP-I $\uparrow$ DINO $\uparrow$ FGIS $\uparrow$
$\checkmark$ - 66.4 77.8 82.9 - 75.5 86.1 85.6
$\checkmark$ 75.5 86.1 85.6 $\checkmark$ 71.0 86.5 85.7
Facial Feature CLIP-I $\uparrow$ DINO $\uparrow$ FGIS $\uparrow$ ImageProjection CLIP-I $\uparrow$ DINO $\uparrow$ FGIS $\uparrow$
Overall Facial Feature 72.9 80.7 84.2 - 61.0 75.6 82.9
Fine-grained Feature 75.2 86.6 85.4 $\checkmark$ 75.5 86.1 85.6
Overall Facial & Fine-grained Feature 75.5 86.1 85.6
Table 2: Ablation study on ID features, loss functions, ImageProjection module, and the usage of LLaVA1.5 in inference.
Facial attention localization strategy: We investigated the effectiveness of facial attention localization strategies during training. The first strategy involves $\mathcal{L}_{noise}$, while the second strategy adds attention loss $\mathcal{L}_{loc}$. From Table 2, we observe that ConsistentID experiences a clear improvement in metrics related to facial feature consistency and fine-grained ID preservation when $\mathcal{L}_{loc}$ is considered. This confirms the effectiveness of maintaining ID consistency between facial regions and the entire face during the training process.
Image projection module: Additionally, we compared two training strategies. The first involves frozen weights of the image projection model and only training our designed FacialEncoder. The second strategy involves training both simultaneously. The results from Table 2 indicate that concurrently training ImageProjection brings the maximum benefits to the model. This is attributed to our model being an ID preservation method of cooperative training with multimodal text and image information.
The LLaVA1.5 usage: In Table 2 (bottom), we present comparative results conducted with textual descriptions using and not using LLaVA 1.5. When LLaVA1.5 is not utilized, we use the prompt ‘The person has one nose, two eyes, two ears, and a mouth’ to replace detailed descriptions of facial regions. We observe an improvement in the generated quality with the introduction of LLaVA1.5 in most metrics, with only CLIP-I showing degradation. This degradation is attributed to the local attention of CLIP, which tends to overlook finer facial details.
4.4 Applications
Following the approach proposed in Photomaker [22], we verify facial high-fidelity and naturalness through two downstream applications presented in Figure 10. These applications involve ‘Change Age & Gender’ and ‘Bringing a person in an old photo into reality’. Compared to Photomaker and InstantID, our ConsistentID demonstrates robust capabilities in maintaining ID consistency, facial high-fidelity, and natural expressions. This superiority is attributed to the effective fine-grained ID preservation of facial features.
Figure 10: The comparisons of two downstream applications.
5 Conclusion and Limitation
Limitations: The utilization of MLLM in our approach may introduce limitations that could affect specific facets of model performance. The constraints posed by limited pose and expression may restrict the diversity of our method, impacting its capability to handle facial variations. These limitations underscore the necessity for in-depth discussion and exploration, specifically in addressing challenges related to pose, expression, and the integration of GPT-4V.
Appendix Materials of ConsistentID
Jiehui Huang Xiao Dong Wenhui Song Hanhui Li Jun Zhou Yuhao Cheng Shutao Liao Long Chen Yiqiang Yan Shengcai Liao Xiaodan Liang
6 Compare with more fine-tuning-based models
We individually fine-tune CustomDiffusion [20], Dreambooth[40], and LoRA [12] using 192 face images of Obama and 158 of Taylor Swift. These fine-tuned models are then utilized for text-to-image inference. To ensure a fair comparison, all models are based on Stable Diffusion v1.5 [37]. In Figure 11, we compare the proposed ConsistentID with these fine-tuned models. It is noteworthy that our training of ConsistentID relies solely on a single reference image, yet it achieves comparable quality in synthesized images to methods trained with hundreds of reference images.
Figure 11: The comparisons with more fine-tuning-based models.
7 More ablation
Attention loss $\mathcal{L}_{loc}$: To further validate the effectiveness of our $\mathcal{L}_{loc}$, we present several comparisons using two different identities in Figure 12. From the figure, we draw the following conclusions: 1) The details of key facial areas, such as Li Feifei’s eye shape, are well preserved. 2) The appearances of facial regions are highly consistent with the reference images, like eyes, ears, and mouths.
Figure 12: Visualized results with or without using attention loss.
Delay control: In Figure 13, we provide visualized results to evaluate the impact of delay control during inference. The term ‘merge step’ denotes the first time step in which we incorporate fine-grained facial image features. It helps to control the balance between text prompts and face images. In general, as the ‘merge step’ increases, the influence of fine-grained facial image features gradually diminishes. For example, if the ‘merge step’ is set to 0, it indicates that fine-grained facial image features are dominant in the generation process and might result in semantic inconsistencies. On the contrary, setting the ‘merge step’ to 0 will maximize the guidance of text prompts, yet might harm the identity consistency.
Infer time of each module: In Table 3, we present the inference time of each module, which adds up to 16 seconds for processing one image. It quantitatively demonstrates the competitive performance in terms of inference efficiency  [^2] During the inference phase, the LLaVa or ChatGPT--4v module is not necessary input. Effective outcomes can be achieved by solely utilizing predefined descriptors of the facial description ‘face, nose, eyes, ears, and mouth’..
BiSeNet InsightFace FacialEncoder UNet Inference
Time (s) 1 3 3 5 4
Table 3: Inference time of each module.
8 Dataset details
@@
Data source: Our facial images are from three public datasets, including FFHQ [17], CelebA[27], and SFHQ[2]. We select 70,000, 30,000, and 424,258 images from these datasets, respectively. In a total of 524,258 images, 107,048 images have recognizable IDs. Figure 16 shows some examples of these images and their corresponding captions.
Dataset pipeline: Below is the detailed our dataset processing pipeline: Each image is initially fed into BiSeNet and LLaVA1.5 to obtain a fine-grained mask and facial descriptions. Let $I_{\text{mask}}$ denote the fine-grained image mask, which is a binary mask indicating our defined facial components. We then utilize $I_{\text{mask}}$ to segment out the corresponding regions from the original image and denote them as $I_{\text{face}}$. Meanwhile, we utilize the InsightFace model to extract facial identity features.
Figure 16: Several training data demos from our FGID dataset.
@@
[…]
Table 4: ID names used for evaluation.
Dataset characteristics: In our dataset, we include 15 types of identities, which are listed in Table 4. Subsequently, in Figure  17, we display the distributions of gender and age across all identities. The figure illustrates a relative balance of both properties in our dataset.
Dataset scenarios: In table 5, we further provide all prompts used in 45 scenarios. The scenarios are divided into 4 categories based on the different applications, including Clothing&\&&Accessory, Action, Background and Style.
@@
Table 5: Evaluation text prompts are categorized by Clothing&Accessories, Background, Action, and Style. During inference, the term ‘class’ will be substituted with ‘man’, ‘woman’, ‘girl’, etc. For each identity and prompt, we generated 1,000 images randomly for evaluation.
9 More training details.
1. We define the keywords of $T_{\text{face}}$ as ‘face’, ‘ears’, ‘eyes’, ‘nose’, and ‘mouth’. These keywords are then used to locate their positions in $I_{\text{face}}$, and a trigger word ‘<facial>’ is inserted to replace the keywords at the matched positions. Next, the descriptions corresponding to facial regions are rearranged according to the order of facial feature keywords, while any descriptions unrelated to facial regions are eliminated. To resolve the problem of incomplete alignment between facial region descriptions and actual facial regions, we address two scenarios. If $T_{\text{face}}$ comprises fewer than 5 region descriptions, $I_{\text{face}}$ retains only the corresponding regions, with unidentified regions substituted by zero matrices. Conversely, if $T_{\text{face}}$ includes complete descriptions for 5 regions but $I_{\text{face}}$ lacks descriptions for all 5, the absent region is replaced with a zero matrix.
2. $T_{\text{face}}$ with trigger words undergoes ID encoding via a tokenizer, resulting in an encoded vector $input\_ids$ of length 77. The positions of trigger word representations in $input\_ids$ are recorded for FacialEncoder model localization during trigger word replacement.
3. To extract prior information for overall face ID, we referred to the IP-Adapter model and employed the InsightFace model to extract facial ID features for all images.
10 More visualization
Downstream applications: In Figures 17, 18, 19, and 20, we present additional visualized results to demonstrate the capabilities of our model in high-fidelity and flexible editing across various recontextualization scenarios, including age modification, identity mixing, and gender transformation.
Figure 17: Additional application cases of ConsistentID for altering the age attribute of a character.
Figure 18: Additional application cases of ConsistentID for identity confusion. Utilizing the overall facial ID feature of one character (top). Leveraging the fine-grained multimodal features from another character (bottom).
Figure 19: Additional application cases of ConsistentID for bringing old photos back to life.
Figure 20: Extended visualization in re-contextualization settings. These examples demonstrate the high-identity fidelity and text editing capability of ConsistentID.
11 Discussion of ethical principles:
In the work, we introduce ConsistentID, a method for generating high-quality facial images while preserving identity fidelity. Our approach emphasizes efficiency, diversity, and controllability in facial generation tasks, serving as a robust baseline for academic research. However, the widespread adoption of such technology raises ethical concerns regarding privacy, misinformation, and potential misuse. We advocate for the responsible development and use of these tools, emphasizing the importance of ethical guidelines to ensure their safe and ethical application in computer vision.

### v2

@@ 3.3 Training and Inference Details
The training data for ConsistentID consists of facial image-text pairs. In ConsistentID, to enhance text controllability, we prioritize the caption as the primary prompt and concatenate it with more detailed descriptions of facial regions extracted from LLaVA1.5, forming the ultimate textual input. During training, only the parameters of the facial encoder and projection module within the facial ID feature extractor are optimized, while the pre-trained diffusion model remains frozen. Regarding training loss functions, they align with those used in the original stable diffusion models and are expressed as:
@@
The total loss function is $\mathcal{L}_{total}=\mathcal{L}_{noise}+\lambda\mathcal{L}_{\text{facial}}$.
During the inference process, we employ a straightforward delayed primacy condition as similar to Fastcomposer. This allows the use of a separate text representation initially, followed by enhanced text representation after a specific step, effectively balancing identity preservation and editability.
3.4 FGID Dataset
@@
To address this limitation, we introduce the FGID dataset, which provides comprehensive fine-grained ID information and detailed facial descriptions essential for training the ConsistentID model. In the following, we will describe the dataset curation process and key characteristics of this dataset.
3.4.1 Dataset Curation
@@
Data source: Our facial images in the FGID dataset are from three public datasets, including FFHQ [40], CelebA [26], and SFHQ [54]. We separately select 70,000, 30,000, and 424,258 images from these datasets. Finally, a total of 524,258 images are selected, where 107,048 images have recognizable IDs. Figure 6 shows some examples of these images and their corresponding captions.
Dataset Pipeline: As for the selected images, each image is first processed through BiSeNet [50] to generate a fine-grained binary mask ($I_{\text{mask}}$) that identifies predefined facial components. Using $I_{\text{mask}}$, the corresponding facial regions are segmented from the original image, denoted as $I_{\text{face}}$. Meanwhile, the InsightFace [56] model is used to extract comprehensive facial identity features, ensuring both global and regional identity information is captured effectively.
For textual data, the LLaVA1.5 model generates detailed descriptions using the embedded prompt: ‘Please describe the people in the image, including their gender, age, clothing, facial expressions, and any other distinguishing features.’. This ensures fine-grained textual annotations are paired with the visual data.
3.4.2 Dataset Characteristics
The FGID dataset is designed to provide comprehensive textual and visual information for training models like ConsistentID. Below, we summarize its key features and characteristics:
Dataset Composition: The FGID dataset includes 15 distinct identity types, as detailed in Table I. To ensure diversity and inclusiveness, the distributions of gender and age across all identities are presented in Figure 4, demonstrating a relatively balanced representation of these properties.
Dataset Scenarios: The dataset spans 45 scenarios, categorized into four application areas: Clothing & Accessory, Action, Background, and Style. Table II lists the prompts used to define these scenarios, ensuring a wide range of contextual diversity.
Rich Data Representation: The FGID dataset is tailored to capture both whole-face and specific facial feature information, providing fine-grained textual and visual details essential for model training. This design enhances its utility for generating accurate and diverse facial representations.
Future Directions: To further augment diversity and information content, we plan to expand the dataset by incorporating additional resources, such as higher-level datasets like LAION-Face [24] and self-collected multi-ID data.
@@ 3.4.2 Dataset Characteristics
1. Andrew Ng 6. Scarlett Johansson 11. Joe Biden
2. Barack Obama 7. Taylor Swift 12. Kamala Harris
3. Dwayne Johnson 8 Albert Einstein 13. Kaming He
4. Fei-Fei Li 9. Elon Mask 14. Lecun Yann
5. Michelle Obama 10. Geoffrey Hinton 15. Sam Altman
TABLE I: ID names used for evaluation.
@@
TABLE II: Evaluation text prompts are categorized by Clothing&Accessories, Background, Action, and Style. During inference, the term ‘class’ will be substituted with ‘man’, ‘woman‘, ‘girl‘, etc. For each identity, we use these prompts to generate 45 images for evaluation.
4 Experiments
Figure 5: User preferences across image fidelity, fine-grained ID fidelity, overall ID fidelity for different methods.
Figure 6: Several training data demos from our FGID dataset.
4.1 Implementation Details
Experimental Implements. In ConsistentID, we employ the Stable Diffusion V1.5 (SD1.5) as the foundational text-to-image training model. For the fine-grained multimodal feature extractor, we initialize the parameters of all text encoders and image encoders with CLIP-ViT-H  [57]. Additionally, we use the image projection layers from CLIP-ViT-H to initialize the learnable projection module within the overall facial ID feature extractor. The entire framework is optimized using Adam [58] on 8 NVIDIA 3090 GPUs, with a batch size of 16. We set the learning rate for all trainable modules to $1\times 10^{-4}$. During training, we probabilistically remove 50% of the background information from the characters with a 50% probability to mitigate interference. The hyperparameter coefficient ($\lambda$) of facial features loss $\mathcal{L}_{\text{facial}}$ is set to 0.01. Additionally, to enhance generation performance through classifier guidance, there is a 10% chance of replacing the original updated text embedding with a zero text embedding. During inference, we use delayed topic conditioning  [40, 20] to resolve text and ID condition conflicts. We generate portrait images using a 50-step DDIM sampler [6] with a classifier guidance scale set to 5.
Experimental Metrics. To evaluate the effectiveness and efficiency of ConsistentID, we employ six widely used metrics [34]: CLIP-I [18], CLIP-T [29], DINO [59], FaceSim [60], FID [61], and inference speed. CLIP-T measures the average cosine similarity between prompt and image CLIP embeddings, evaluating ID fidelity. CLIP-I calculates the average pairwise cosine similarity between CLIP embeddings of generated and real images, assessing prompt fidelity. DINO represents the average cosine similarity between ViT-S/16 [62] embeddings of generated and real images, indicating fine-grained image-level ID quality. FaceSim determines facial similarity between generated and real images using FaceNet [60]. FID gauges the quality of the generated images [61]. Inference speed denotes the calculation time under the same running environment. To fully evaluate and quantify quality at the facial region level, we propose a novel metric, FGIS (Fine-Grained Identity Similarity). FGIS is computed as the average cosine similarity between DINO embeddings of the generated facial regions in reference and generated images. A higher FGIS value indicates greater ID fidelity in the generated facial regions.
CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINO $\uparrow$ FaceSim $\uparrow$ FGIS $\uparrow$ FID $\downarrow$ Speed (s)
Fastcomposer [20] 27.8 67.0 68.4 75.2 77.7 372.8 10
IP-Adapter [15] 27.6 75.0 74.5 75.6 73.4 320.0 13
Photomaker [21] 30.7 71.7 72.6 69.3 73.2 336.5 17
InstantID [22] 30.3 68.2 77.6 76.5 78.3 271.9 19
ConsistentID 31.1 76.7 78.5 77.2 81.4 312.4 16
TABLE III: Quantitative comparison of the universal recontextualization setting on the MyStyle test dataset. The benchmark metrics assessed text consistency (CLIP-T), the preservation of coarse- and fine-grained ID information (CLIP-I, DINO, FaceSIM, and FGIS), generation quality (FID), and inference efficiency (speed in seconds).
4.2 Comparation Results
To demonstrate the effectiveness of ConsistentID, we conduct a comparative analysis against state-of-the-art methods, including Fastcomposer [20], IP-Adapter[15], Photomaker [21], and InstantID [22]. During testing, we utilized the officially provided models and generate portrait images using only a single reference image. Consistent with the evaluation protocol of Photomaker, we utilize the Mystyle [27] dataset for quantitative assessment and incorporate over ten identity datasets for qualitative visualization.
Figure 8: Qualitative comparison of our model with other models on the task of decoupling ID and text instructions. Notably, our method can control both the expression and style of the ID, and also demonstrates excellent results in maintaining the identity of anime characters.
Quantitative results: Following Photomaker [21], we used the test dataset from Mystyle [27]. Notably, during the inference process, we did not use LLaVA to generate the descriptions of each facial region but obtained facial feature information using a fixed phrase: ‘This person has one nose, two eyes, two ears, and a mouth.’ as the text prompt. The quantitative comparison is conducted under the universal recontextualization setting, utilizing a set of metrics to benchmark various aspects.
The results are showcased in Table III. A thorough analysis of the table demonstrates that ConsistentID consistently outperforms other methods across most evaluated metrics, and surpasses other IP-Adapter-based methods in terms of generation efficiency. This is attributed to ConsistentID’s fine-grained ID preservation capability and the efficiency of the lightweight multimodal facial prompt generator. Regarding the FID metric, the relatively lower performance may be influenced by the inherent generative limitations of the base model SD1.5.
Visualized comparisons under different scenarios: To demonstrate the advantages of ConsistentID visually, we present the text-edited generation results of all methods, using reference images from five distinct identities, in Figure 7. This visualization highlights ConsistentID’s capability to produce vibrant and realistic images, with a particular emphasis on facial features. To further elucidate this observation, we selectively magnify and compare specific facial details across all methods in four identities, as depicted in Figure 9. Our model showcases exceptional ID preservation capabilities in facial details, especially in the eyes and nose, attributed to fine-grained multimodal prompts and facial regions’ ID information.
Figure 10: The comparisons with more fine-tuning-based and direct inference models. For direct inference method GPT4o, we use stars’ names in the text prompt.
To validate the capability of accurate text understanding, we additionally present style-based and action-based text-edited results in Figure 8. It can be observed that the images generated by InstantID exhibit limited flexibility in facial poses. This limitation is likely attributed to Controlnet-based prompt insertion methods, which may easily overlook textual prompts. Simultaneously, we notice that, while Photomaker can accurately comprehend textual and visual prompts, it lacks the ID consistency of facial regions. In contrast, our ConsistentID achieves optimal generation results due to its precise understanding of textual and visual prompts. This further emphasizes the importances of multimodal fine-grained ID information. To fully show the advantages of our ConsistentID, more visualized comparisons are provided in following section, including comparative experiments with fine-tuning-based models (Facechain [63], and Easyphoto [64]), and the MLLM model GPT4o.
Figure 11: Results of our proposed ConsistentID model extended on the SDXL base model, compared with other SDXL-based models.
4.3 Compare with more models
To further illustrate the advantages of our model, we compare ConsistentID with two LoRA-based fine-tuning methods, Facechain [63] and Easyphoto [64], as well as the closed-source model GPT4o  [^2] https://chatgpt.com/. To ensure a fair comparison, all models are provided with a single reference image as input. For the LoRA-based methods, the input comprises image + prompt, where image represents the reference image and prompt specifies the editing instruction for generating personalized outputs. For the GPT4o model, the template used is: ‘Edit this image using the following instruction: person_name + prompt’, where person_name specifies the individual’s name. In Figure 10, we show the visualized results of the personalization comparison using a single image, demonstrating the superior capability of our ConsistentID in keeping ID consistency.
In addition, we compare ConsistentID with models specifically designed using IP-Adapter as the base model, as shown in Figure 12. The figure clearly demonstrates that these models struggle to achieve detailed ID preservation in facial regions, primarily due to the lack of fine-grained textual and visual prompts. In contrast, ConsistentID exhibits a robust capability to preserve the integrity of facial ID and seamlessly blend it into various styles, utilizing multimodal fine-grained prompts. This highlights ConsistentID’s advantage in retaining identity while providing enhanced flexibility and control over style customization.
4.4 Human Study
To accurately capture user preferences, we conducted surveys to evaluate perceptions of image fidelity, fine-grained ID fidelity, and overall ID fidelity. Figure 5 visualizes the proportion of total votes received by each method. Across all three metric dimensions, ConsistentID achieves the highest user preference share. This result highlights the effectiveness of ConsistentID in meeting user expectations for both identity preservation and image quality.
4.5 Ablation Study
Facial ID Types: We conducted an ablation study on facial IDs, considering three variations: using only the holistic facial ID, using our designed fine-grained IDs, and using both the holistic facial ID and our designed fine-grained IDs. The results in Table IV indicate that using fine-grained ID features alone effectively maintains ID consistency in the facial region, particularly evidenced by an increase in DINO values. In contrast, using only overall facial features (FaceID) results in decreased values across all metrics, due to the coarse-grained characteristic of the overall facial feature. Finally, the combination of holistic facial ID and fine-grained ID features achieves the best overall ID fidelity, demonstrating the complementary strengths of both approaches and their contribution to enhanced model performance.
$\mathcal{L}_{noise}$ $\mathcal{L}_{facial}$ CLIP-I $\uparrow$ DINO $\uparrow$ FGIS $\uparrow$ LLaVA1.5 CLIP-I $\uparrow$ DINO $\uparrow$ FGIS $\uparrow$
-- 66.4 77.8 82.9 -- 67.4 84.3 84.9
75.5 86.1 85.6 75.3 85.4 85.8
Facial Feature CLIP--I $\uparrow$ DINO $\uparrow$ FGIS $\uparrow$ ImageProjection CLIP--I $\uparrow$ DINO $\uparrow$ FGIS $\uparrow$
Overall Facial Feature 72.9 80.7 84.2 -- 61.0 75.6 82.9
Fine-grained Feature 75.2 86.6 85.4 75.5 86.1 85.6
Overall Facial & Fine-grained Feature 75.5 86.1 85.6
TABLE IV: Ablation study on ID features, loss functions, ImageProjection module, and the usage of LLaVA1.5 in training.
Figure 13: Ablation study of the number of facial regions.
Figure 14: Visualized results with or without using attention loss.
Facial attention localization strategy: We investigated the effectiveness of facial attention localization strategies during training. The first strategy involves $\mathcal{L}_{noise}$, while the second strategy adds attention loss $\mathcal{L}_{facial}$. From Table IV, we observe that ConsistentID experiences a clear improvement in metrics related to facial feature consistency and fine-grained ID preservation when $\mathcal{L}_{facial}$ is considered. This confirms the effectiveness of maintaining ID consistency between facial regions and the entire face during the training process.
Image projection module: Additionally, we compared two training strategies. The first involves frozen weights of the image projection model and only training our designed FacialEncoder. The second strategy involves training both simultaneously. The results from Table IV indicate that concurrently training ImageProjection brings the maximum benefits to the model. This is attributed to our model being an ID preservation method of cooperative training with multimodal text and image information.
The LLaVA1.5 usage: In Table IV (bottom), we analyze the impact of using LLaVA 1.5 to generate textual descriptions during training. Our observations indicate that incorporating LLaVA 1.5 leads to noticeable improvements across various metrics, with particularly significant gains in the CLIP-I metric. Notably, during inference, our model achieves satisfactory performance without relying on LLaVA 1.5, using only the simplified facial description, ‘This person has one nose, two eyes, two ears, and a mouth,’ instead of generating detailed facial region descriptions. This demonstrates that our model effectively learns fine-grained features during the training phase with LLaVA 1.5, enabling it to maintain strong image understanding and identity preservation during inference.
Attention loss $\mathcal{L}_{facial}$: To further validate the effectiveness of our $\mathcal{L}_{facial}$, we present several comparisons using two different identities in Figure 14. From the figure, we draw the following conclusions: 1) The details of key facial areas, such as Taylor’s eye shape, are well preserved. 2) The appearances of facial regions, such as eyes, ears, and mouths, effectively reflect identity consistency with the reference images.
Additionally, as shown in Figure 17, the attention scores in the model’s facial feature regions gradually increase as training progresses. This indicates improved semantic alignment between the facial regions and their corresponding textual descriptions, further supporting the efficacy of $\mathcal{L}_{facial}$.
Delay control: In Figure 15, we provide visualized results to evaluate the impact of delay control during inference. The term ‘merge step’ denotes the first time step in which we incorporate fine-grained facial image features. It helps to control the balance between text prompts and face images. In general, as the ‘merge step’ increases, the influence of fine-grained facial image features gradually diminishes.
For example, if the ‘merge step’ is set to 0, it indicates that fine-grained facial image features are dominant in the generation process and might result in semantic inconsistencies. On the contrary, setting the ‘merge step’ to 0 will maximize the guidance of text prompts, yet might harm the identity consistency.
Inference time for each module: In Table V, we show the processing time for each module involved in handling a single image during inference, with a total duration of 16 seconds per image. This quantitative analysis highlights the competitive performance of the system in terms of inference efficiency  [^3] Notably, during the inference phase, the usage of the LLaVa or ChatGPT-4v module is not mandatory. Satisfactory results can be obtained by relying exclusively on predefined facial features prompts, such as ‘face, nose, eyes, ears, and mouth.’.
Extensibility of methods for base models: We further replace the base model in the ConsistentID framework with the SDXL model, as illustrated in Figure 11. From the figure, ConsistentID effectively enhances facial feature consistency and outperforms other models in terms of ID similarity. Furthermore, we conducted additional ablation experiments using the SDXL model as the base for the ConsistentID framework, as shown in Table VI. Compared to the SD1.5-based ConsistentID model, the SDXL-based version demonstrates notable improvements across key metrics, including facial similarity, CLIP-I, CLIP-T, and FID. These results further validate the robust generalization capabilities of the ConsistentID model and demonstrate its flexibility in utilizing more advanced base models for improved performance.
BiSeNet InsightFace FacialEncoder UNet Inference
Time (s) 1 3 3 5 4
TABLE V: Inference time of each module.
CLIP-T $\uparrow$ FaceSim $\uparrow$ FID $\downarrow$ Speed (s) $\downarrow$
ConsistentID-SD1.5 19.8 62.9 307.9 16
ConsistentID-SDXL 21.9 73.1 304.5 18
TABLE VI: ConsistentID extension ablation experiment, evaluating the performance using different base models. The average performance is tested on 180 images under 45 prompt conditions, using the character ID shown in Figure 11.
Figure 17: Experiments to visualize facial changes in attention map during training results.
5 Conclusion